### v1

# Direct Adaptive Control of Grid-Connected Power Converters via Output-Feedback Data-Enabled Policy Optimization

Abstract
Power electronic converters are gradually becoming the main components of modern power systems due to the increasing integration of renewable energy sources. However, power converters may become unstable when interacting with the complex and time-varying power grid. To deal with this problem, an adaptive control design scheme for power converters is preferable, which can capture the closed-loop dynamical interaction between the converter and the grid via online data. In this paper, we propose an adaptive data-driven control method, called data-enabled policy optimization (DeePO), to stabilize power converters by using only online input-output data. Our contributions are threefold. First, we propose a covariance parameterization of partially observed linear systems with input-output data. Second, we develop a DeePO algorithm, which updates the parameterized policy with data-based gradient descent to achieve computationally efficient adaptive control. Third, we use high-fidelity simulations to verify that DeePO can effectively stabilize grid-connected power converters and quickly adapt to the changes in the power grid.
I Introduction
Modern power systems feature a large-scale integration of power electronic converters, as they act as interfaces between the AC power grid and renewable energy sources, high-voltage DC (HVDC) systems, energy storage systems, and electric vehicles [1]. The large-scale integration of converters is fundamentally changing the power system dynamics, as they are significantly different from traditional synchronous generators (SGs). Usually, multiple nested control loops, based on fixed-parameter PI regulators, are needed in converters to achieve voltage, current, and power regulations. Under these control loops, power converters exhibit complicated interaction with the power grid and may easily tend to be unstable due to unforeseen grid conditions [2, 3, 4]. Such instability issues have been widely observed in practice [5], which poses challenges to the secure operation of modern power systems and impedes further integration of renewables.
The instability in converter systems is caused by the closed-loop interaction between the converter and the complex power grid, which often occurs when the converter’s control strategy does not fit into the grid characteristics [6, 7]. Hence, the control design of converters should take into account the power grid dynamics for the sake of stability. However, the power grid is unknown, nonlinear, and time-varying from the perspective of a converter. Moreover, the grid structure and parameters are difficult to obtain in real time. Hence, it is nearly impossible to establish an exact dynamical model of a power grid for the control design of converters. As a remedy, engineers often use an overly simplified model for the power grid (e.g., an infinite bus) and tune the controller based on engineering experience and iterative trial-and-error approaches, which can be expensive, time-consuming, and lacking stability guarantees due to the model mismatch. While existing robust control methods can be used to handle the model mismatch [8, 9], they usually lead to conservative controllers when large changes appear in the power grid (e.g., tripping of transmission lines or even HVDC stations). Ideally, the controller of converters should be adaptive, i.e., it is able to perceive and quickly adapt to changes in the power grid by using online data.
Recently, there has been a renewed interest in direct data-driven control, which bypasses the system identification (SysID) step and learns the controller directly from a batch of persistently exciting data [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. This approach is end-to-end, easy to implement, and has seen many successful applications [20, 21, 22]. Following this line, our previous work proposes a direct data-driven linear quadratic regulator (LQR) method [23, 24]. It is adaptive in the sense that the control performance is improved in real time by using online closed-loop data. We call this method Data-enabled Policy Optimization (DeePO), where the policy is parameterized with sample covariance of input-state data and updated using gradient methods. DeePO has a recursive algorithmic implementation and hence is computationally efficient. From a theoretical perspective, it meets provable stability and convergence guarantees for linear time-invariant systems. All these key features render DeePO an ideal data-driven control method for power converters.
In this paper, we propose a DeePO method to mitigate oscillations in power converter systems. Since the power converter system is only partially observed, i.e., the full state is unmeasurable, we first extend the covariance parameterization of the system with input-state data in [23, 24] to that with input-output data. This is achieved by regarding a finite length of past input-output trajectory as the state and using the corresponding covariance to parameterize the closed-loop system. Then, we propose the DeePO algorithm for direct adaptive control of partially observed systems. Finally, we apply the DeePO algorithm to stabilize power converter and direct-drive wind generator systems, both of which are unknown, state-unmeasurable, and encounter sudden change of dynamics due to grid changes. Simulation results show that DeePO enables efficient online adaptation and effectively prevents instabilities in grid-connected power converters. Compared with our previous works applying data-enabled predictive control (DeePC) to power converters [21, 25], DeePO offers a significantly reduced online computational burden, making it more suitable for scenarios where the processor cannot solve a quadratic program in real time.
The rest of the paper is organized as follows. Section II formulates the stabilization problem of power converter systems. Section III introduces the adaptive LQR control and the DeePO method. Section IV extends the DeePO algorithm to input-output systems. Section V performs simulations on the converter systems. Conclusion is made in Section VI.
Notation. We use $I_{n}$ to denote the $n$-by-$n$ identity matrix. We use $\rho(\cdot)$ to denote the spectral radius of a square matrix. We use $A^{\dagger}$ to denote the pseudoinverse of a matrix $A$. We use $(S)_{i}$ to denote the $i$-th column of a block matrix $S$.
II Stabilization of grid-connected power converters
In this section, we introduce the stabilization problem of DC-AC power converter systems, taking into account both the DC-side and AC-side dynamics. Notice that power converters are widely used as the interface between the AC power grid and a DC source, such as lithium battery-based energy storage systems or direct-drive wind generators. The lithium batteries generate a constant DC voltage, and the converter aims to regulate the active and reactive power, as shown in Fig. 1. By comparison, when used as the grid interface of direct-drive wind generators, the converter needs to regulate the DC voltage and the reactive power, as will be shown in Section V. In both cases, the converter plays an important role in maintaining stable and reliable power transfer between the DC side and the AC side.
II-A Power converter systems
Figure 1: One-line diagram of a grid-connected power converter. Here the DC side is connected to lithium batteries, while it can also be wind turbines.
Consider the grid-connected power converter system in Fig. 1, which has a phase-locked loop (PLL), power/current control loops, and (abc/dq) coordinate transformation blocks [6]. Due to proprietary manufacturer models and the complexity of the power grid, the power converter together with the grid is a black-box system for the subsequent stabilization control design. While inherently nonlinear, the system can be linearized around its equilibrium point for analysis. Note that the time-varying nature of the power grid may induce a time-varying operating point. Without loss of generality, consider the state-space model linearized at the origin
$x_{t+1}$ $=Ax_{t}+Bu_{t}+w_{t},$ (1)
$y_{t}$ $=Cx_{t}+v_{t},$
where $x_{t}\in\mathbb{R}^{n}$ is the state variable (could be unmeasurable, e.g., variables in the power grid side), $u_{t}\in\mathbb{R}^{m}$ is the control input (e.g., additional signals added to the current references; see Fig. 1), $y_{t}\in\mathbb{R}^{d}$ is the output (e.g., active and reactive power in our setting), $w_{t}$ is the process noise, and $v_{t}$ is the measurement noise. The unknown system $(A,B,C)$ is controllable and observable, but it may be subject to sudden impedance changes due to external grid changes, e.g., changes of $Z_{\rm grid}$ in Fig. 1. Such events may lead to poorly damped or even unstable oscillations.
Our objective is to find a policy as feedback of the past input-output trajectory $u_{t}=\pi_{t}(u_{-\infty},y_{-\infty},\dots,u_{t-1},y_{t-1})$ so that the output is regulated to zero with low control effort.
II-B Challenges in stabilization of power converter systems (DC source + DC-AC converter + AC power grid)
There are three main challenges in the stabilization of power converter systems:
  • unknown model: Power converter systems often exhibit high-order characteristics, rendering it difficult to establish exact dynamical models. Specifically, the exact dynamics of the DC source and the converter are usually hard to obtain (and therefore unknown) due to proprietary manufacturer models and their complexity, e.g., the complex mechanical behaviors of wind turbines. Moreover, the power grid is also unknown when designing a controller for the converter, as the grid is large-scale and time-varying (therefore hard to model).
  • measurement limitations: Internal states are not measurable and hence we can only use input-output information for the control design.
  • sudden change of dynamics: Variations in grid conditions, operating states of power converters, and control parameter adjustments cause the power converter system to encounter sudden changes in dynamics [21].
II-C Our approach
In this paper, we propose a direct data-driven method to solve the stabilization problem of power converter systems without SysID. Our approach is based on data-enabled policy optimization (DeePO), an adaptive linear quadratic control method introduced in our recent work [23, 24]. It is data-driven and does not involve any explicit SysID. Moreover, DeePO uses online closed-loop data to adaptively update the control policy, making it an ideal method to deal with time-varying dynamics.
In the sequel, we first recapitulate the DeePO method with input-state data in [23, 24]. Then, we extend DeePO to the input-output setting. Finally, we discuss its implementation on power converter systems and perform simulations to validate the effectiveness of DeePO.
III Direct adaptive LQR control via data-enabled policy optimization
In this section, we first introduce the linear quadratic regulator (LQR). Then, we recapitulate data-enabled policy optimization for adaptive learning of the LQR [23, 24].
III-A The linear quadratic regulator
Consider a linear time-invariant system with fully measurable state
[…]
Here, $h_{t}$ is the performance signal of interest, $(A,B)$ is controllable, and the weighting matrices $(Q,R)$ are positive definite.
The LQR problem is phrased as finding a state-feedback gain $K\in\mathbb{R}^{m\times n}$ that minimizes the $\mathcal{H}_{2}$-norm of the transfer function $\mathscr{T}(K):w\rightarrow h$ of the closed-loop system
$\begin{bmatrix}x_{t+1}\\ h_{t}\end{bmatrix}=\begin{bmatrix}A+BK&I_{n}\\ \hline\cr\begin{bmatrix}Q^{1/2}\\ R^{1/2}K\end{bmatrix}&0\end{bmatrix}\begin{bmatrix}x_{t}\\ w_{t}\end{bmatrix}.$
When $A+BK$ is stable, it holds that [26]
$\|\mathscr{T}(K)\|_{2}^{2}=\text{Tr}((Q+K^{\top}RK)\Sigma_{K})=:C(K),$ (3)
where $\Sigma_{K}$ is the closed-loop state covariance matrix obtained as the positive definite solution to the Lyapunov equation
$\Sigma_{K}=I_{n}+(A+BK)\Sigma_{K}(A+BK)^{\top}.$ (4)
We refer to $C(K)$ as the LQR cost and to (3)-(4) as a policy parameterization of the LQR.
It is well-known that the optimal LQR gain $K^{*}$ is unique and can be found by, e.g., solving an algebraic Riccati equation with $(A,B)$ [26]. When $(A,B)$ is unknown, data-driven methods learn the LQR gain from input-state data. In the sequel, we recall the DeePO approach proposed in our previous work [23, 24].
III-B DeePO for direct adaptive LQR control
Policy optimization (PO) refers to a class of direct design methods, where the policy is parameterized and recursively updated using gradient methods [27]. In particular, DeePO uses the sample covariance of input-state data to parameterize the policy [24]. Consider the $t$-long time series of states, inputs, noises, and successor states
$X_{0,t}$ $:=\begin{bmatrix}x_{0}&x_{1}&\dots&x_{t-1}\end{bmatrix}\in\mathbb{R}^{n\times t},$
$U_{0,t}$ $:=\begin{bmatrix}u_{0}&u_{1}&\dots&u_{t-1}\end{bmatrix}\in\mathbb{R}^{m\times t},$
$W_{0,t}$ $:=\begin{bmatrix}w_{0}&w_{1}&\dots&w_{t-1}\end{bmatrix}\in\mathbb{R}^{n\times t},$
$X_{1,t}$ $:=\begin{bmatrix}x_{1}&x_{2}&\dots&x_{t}\end{bmatrix}\in\mathbb{R}^{n\times t},$
which satisfy the system dynamics
$X_{1,t}=AX_{0,t}+BU_{0,t}+W_{0,t}.$ (5)
Assuming that the data is persistently exciting (PE) [28], i.e., the block matrix of input and state data $D_{0,t}:=[U_{0,t}^{\top},X_{0,t}^{\top}]^{\top}$ has full row rank
$\text{rank}(D_{0,t})=m+n,$ (6)
the covariance parameterization[24] of $K$ is given by
$\begin{bmatrix}K\\ I_{n}\end{bmatrix}=\frac{1}{t}D_{0,t}D_{0,t}^{\top}V=:\Phi_{t}V,$ (7)
where $\Phi_{t}\succ 0$ denotes the sample covariance of input-state data, and $V\in\mathbb{R}^{(n+m)\times n}$ is the parameterized policy.
With the covariance parameterization (7), the LQR problem (3)-(4) can be expressed by raw data matrices $(X_{0,t},U_{0,t},X_{1,t})$ and the optimization matrix $V$. For brevity, let $\overline{X}_{0,t}=X_{0,t}D_{0,t}^{\top}/t$ and $\overline{U}_{0,t}=U_{0,t}D_{0,t}^{\top}/t$ be a partition of $\Phi_{t}$, and let $\overline{W}_{0,t}=W_{0,t}D_{0,t}^{\top}/t$ be the noise-state-input covariance, and finally define the covariance with respect to the successor state $\overline{X}_{1,t}=X_{1,t}D_{0,t}^{\top}/t$, Note that the dimension of the covariance matrices does not depend on $t$. Then, the closed-loop matrix can be written as
[…]
Following the certainty-equivalence principle [14], we disregard the unmeasurable $\overline{W}_{0,t}$ and use $\overline{X}_{1,t}V$ as the closed-loop matrix. After substituting $A+BK$ with $\overline{X}_{1,t}V$ in (3)-(4) and leveraging (7), the LQR problem becomes
$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V)\Sigma_{t}(V)\right),$ (8)
subject $\text{subject to}~{}~{}\overline{X}_{0,t}V=I_{n},$
where $\Sigma_{t}(V)=I_{n}+\overline{X}_{1,t}V\Sigma_{t}(V)V^{\top}\overline{X}_{1,t}^{\top}$ is a covariance parameterization of (4), and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$. We refer to (8) as the direct data-driven LQR problem, which does not involve any explicit SysID.
The DeePO algorithm uses online gradient descent of (8) to recursively update $V$. The details are presented in Algorithm 1, which alternates between control (line 2) and policy update (line 3-5). The initial stabilizing policy can be either user-specified or computed by solving (8) with offline data. At time $t$, we apply the linear state feedback policy $u_{t}=K_{t}x_{t}+e_{t}$ for control, where $e_{t}$ is a probing noise used to ensure the PE rank condition (6). We use online projected gradient descent in (10) to update the parameterized policy, where the projection $\Pi_{\overline{X}_{0,t+1}}:=I_{n+m}-\overline{X}_{0,t+1}^{\dagger}\overline{X}_{0,t+1}$ onto the nullspace of $\overline{X}_{0,t+1}$ is to ensure the subspace constraint in (8). The gradient can be computed by the following lemma. Define the feasible set of (8) (i.e., the set with stable closed-loop matrices) as $\mathcal{S}_{t}:=\{V\mid\overline{X}_{0,t}V=I_{n},\rho(\overline{X}_{1,t}V)<1\}$.
Lemma 1 ([24])
For $V\in\mathcal{S}_{t}$, the gradient of $J_{t}(V)$ with respect to $V$ is given by
$\nabla J_{t}(V)=2\left(\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}+\overline{X}_{1,t}^{\top}P_{t}\overline{X}_{1,t}\right)V\Sigma_{t}(V),$ (9)
where $P_{t}$ satisfies the Lyapunov equation
$P_{t}=Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V+V^{\top}\overline{X}_{1,t}^{\top}P_{t}\overline{X}_{1,t}V.$
DeePO is direct and adaptive in the sense that it directly uses online closed-loop data to update the policy. Thus, it can rapidly adapt to dynamic changes reflected in the data. Algorithm 1 has a recursive policy update and can be implemented efficiently. Specifically, all covariance matrices and the inverse $\Phi_{t+1}^{-1}$ have recursive updates. Moreover, the parameterization can be updated recursively via rank-one update, i.e.,
$V_{t+1}=\frac{t+1}{t}\left(V_{t}^{\prime}-\frac{\Phi_{t}^{-1}\phi_{t}\phi_{t}^{\top}V_{t}^{\prime}}{t+\phi_{t}^{\top}\Phi_{t}^{-1}\phi_{t}}\right),$
where $\phi_{t}=[u_{t}^{\top},x_{t}^{\top}]^{\top}$, and $\Phi_{t}^{-1}$ and $V_{t}^{\prime}$ are given from the last iteration. Theoretically, it is shown that under mild assumptions the policy $\{K_{t}\}$ converges to the optimal LQR gain. We refer to [24, Section IV] for detailed discussions.
The DeePO framework developed so far considers the state feedback setting. However, the state of power converter system (1) is not measurable, and only input-output data is available for control design. Such an system is typically referred to as partially observed [29]. Next, we extend DeePO to partially observed systems for adaptive control using input-output data.
Algorithm 1 DeePO for direct adaptive LQR control
1:Offline data $(X_{0,t_{0}},U_{0,t_{0}},X_{1,t_{0}})$, an initial policy $K_{t_{0}}$, and a stepsize $\eta$.
2:for $t=t_{0},t_{0}+1,\dots$ do
3:     Apply $u_{t}=K_{t}x_{t}+e_{t}$ and observe $x_{t+1}$.
4:     Update covariance matrices $\Phi_{t+1}$ and $\overline{X}_{1,t+1}$.
5:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
$V_{t+1}=\Phi_{t+1}^{-1}\begin{bmatrix}K_{t}\\ I_{n}\end{bmatrix}.$
6:     Update of the parameterized policy: perform one-step projected gradient descent
$V_{t+1}^{\prime}=V_{t+1}-\eta\Pi_{\overline{X}_{0,t+1}}\nabla J_{t+1}(V_{t+1}),$ (10)
where the gradient $\nabla J_{t+1}(V_{t+1})$ is given by Lemma 1.
7:     Gain update: update the control gain by
$K_{t+1}=\overline{U}_{0,t+1}V_{t+1}^{\prime}.$
8:end for
IV Output-feedback DeePO for partially observed systems
In this section, we extend DeePO for direct adaptive control of partially observed systems using input-output data.
Consider the system (1). Since the state is unmeasurable, we represent (1) with a non-minimal realization using input and output signals. Denote the input trajectory, the output trajectory, and their stack from time $t-n$ to $t-1$ as
${u}_{t,n}=\begin{bmatrix}u_{t-1}\\ \vdots\\ u_{t-n}\end{bmatrix},~{}{y}_{t,n}=\begin{bmatrix}y_{t-1}\\ \vdots\\ y_{t-n}\end{bmatrix},~{}{z}_{t}=\begin{bmatrix}{u}_{t,n}\\ {y}_{t,n}\end{bmatrix}$
respectively. Define the observability matrix
$\mathcal{O}=\begin{bmatrix}CA^{n-1}\\ \vdots\\ CA\\ C\end{bmatrix},$
the controllability matrices ${}\mathcal{C}=[B~{}~{}AB~{}~{}\cdots~{}~{}A^{n-1}B]$, $\mathcal{C}_{w}=[I~{}~{}A~{}~{}\cdots~{}~{}A^{n-1}],$ and the Toeplitz matrices capturing the impulse and disturbance response
$\mathcal{T}=\begin{bmatrix}0&CB&CAB&\cdots&CA^{n-2}B\\ 0&0&CB&\cdots&CA^{n-3}B\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&CB\\ 0&0&0&0&0\end{bmatrix},$
$\mathcal{T}_{w}=\begin{bmatrix}0&C&CA&\cdots&CA^{n-2}\\ 0&0&C&\cdots&CA^{n-3}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&C\\ 0&0&0&0&0\end{bmatrix}.$
Let $S=[C(\mathcal{C}-A^{n}O^{\dagger}\mathcal{T}),CA^{n}O^{\dagger}]$, $d_{t}=C(\mathcal{C}_{w}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}_{w})w_{t,n}-CA^{n}\mathcal{O}^{\dagger}v_{t,n}+v_{t}$. Then, we have the following results, the proof of which is provided in the appendix.
Theorem 1 (Non-minimal controllable realization)
A non-minimal controllable realization of (1) is given by (LABEL:equ:nonnimimal_realization) shown at the bottom of the next page.
Remark 1
The length of the past trajectory used to represent the state does not necessarily match the system order $n$. In fact, it only needs to exceed the system lag (i.e., the observability index), under which the dimension of $z_{t}$ can be reduced.
Then, we can apply DeePO for the system (LABEL:equ:nonnimimal_realization), whose state is measurable. Define the matrix of input-output data
$Z_{0,t}=\begin{bmatrix}z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{n(m+d)\times t}.$
Assume that the data is PE, i.e., the block matrix $[U_{0,t}^{\top},Z_{0,t}^{\top}]^{\top}$ has full row rank. Then, the covariance parameterization of $K$ for the system (LABEL:equ:nonnimimal_realization) is given by
$\begin{bmatrix}K\\ I_{n(m+d)}\end{bmatrix}=\frac{1}{t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top}V=:\Psi_{t}V,$
where $V\in\mathbb{R}^{n(2m+d)\times n(m+d)}$.
Define the sample covariance matrices
$\overline{Z}_{0,t}=\frac{1}{t}Z_{0,t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top},~{}\overline{Z}_{1,t}=\frac{1}{t}Z_{1,t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top}.$
Denote the realization (LABEL:equ:nonnimimal_realization) as $(A_{z},B_{z},C_{z})$. Following the certainty-equivalence principle and disregarding $d_{t}$, we approximate the closed-loop matrix by
$A_{z}+B_{z}K\approx\overline{Z}_{1,t}V.$
Then, the LQR problem of (LABEL:equ:nonnimimal_realization) is parameterized as
$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q_{1}+V^{\top}\overline{U}_{0,t}^{\top}R_{z}\overline{U}_{0,t}V)\Sigma_{t}(V)\right),$ (11)
subject $\text{subject to}~{}~{}\overline{Z}_{0,t}V=I_{n},$
where $Q_{z},R_{z}$ are positive definite matrices with proper dimensions, $\Sigma_{t}(V)=I_{n}+\overline{Z}_{1,t}V\Sigma_{t}(V)V^{\top}\overline{Z}_{1,t}^{\top}$, and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$.
The DeePO method for partially observed systems is provided in Algorithm 2. Next, we apply Algorithm 2 to stabilize power converter and renewable energy systems.
Algorithm 2 DeePO for partially observed systems
1:Offline data $(Z_{0,t_{0}},U_{0,t_{0}},Z_{1,t_{0}})$, an initial policy $K_{t_{0}}$, and a stepsize $\eta$.
2:for $t=t_{0},t_{0}+1,\dots$ do
3:     Apply $u_{t}=K_{t}z_{t}+e_{t}$ and observe $z_{t+1}$.
4:     Update covariance matrices $\Phi_{t+1}$ and $\overline{Z}_{1,t+1}$.
5:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
$V_{t+1}=\Psi_{t+1}^{-1}\begin{bmatrix}K_{t}\\ I_{n}\end{bmatrix}.$
6:     Update of the parameterized policy: perform one-step projected gradient descent
$V_{t+1}^{\prime}=V_{t+1}-\eta\Pi_{\overline{Z}_{0,t+1}}\nabla J_{t+1}(V_{t+1}).$
7:     Gain update: update the control gain by
$K_{t+1}=\overline{U}_{0,t+1}V_{t+1}^{\prime}.$
8:end for
V DeePO for stabilization of power converters and renewable energy systems
In this section, we first discuss the implementation of DeePO on the power converter systems. Then, we perform simulations for a grid-connected power converter and a direct-drive wind generator.
V-A Implementation of DeePO for power converter systems
We consider two typical power converter systems: a grid-connected power converter (with lithium batteries as the DC source) in Fig. 1 and a direct-drive (type 4) wind generator in Fig. 3, which also uses a converter as the grid interface. Both systems share similar control challenges, including unknown models, measurement limitations, and potential change of dynamics. While Algorithm 2 considers only time-invariant systems, it can potentially handle changes in the system dynamics. This is because DeePO uses online closed-loop data to update the policy in real time and can quickly adapt to changes. Moreover, since Algorithm 2 has a recursive implementation, the online update of the policy is computationally efficient.
For both systems, we select the deviations of active and reactive power from their reference values as the outputs of the unknown system, denoted by $y=[P_{E}-P_{0},Q_{E}-Q_{0}]^{\top}$. The DeePO method is configured to provide two control inputs that are added to the current references of the system. All input and output signals are expressed in per-unit ($\mathrm{p.u.}$) values. Since directly measuring internal state variables is challenging, we construct new state variables using input-output information as $z_{t}=[u_{t,T_{\text{ini}}}^{\top},y_{t,T_{\text{ini}}}^{\top}]^{\top}$, where the length of the initial trajectory for constructing new states is set to $T_{\text{ini}}=6$ for both systems.
The parameters for the two systems are set as follows:
  • sampling frequency: The sampling frequency is chosen as $200\,\mathrm{Hz}$.
  • trajectory length $T$: The trajectory length is set to $300$ for the power converter and $600$ for the wind generator to satisfy the persistency of excitation condition.
  • weighting matrices: The input weighting matrix is set as $R=I_{m}$, while the output weighting matrix is set as $Q=\text{diag}(I_{mT_{\text{ini}}},100I_{pT_{\text{ini}}})$.
  • gradient descent parameters: The stepsize for gradient descent is set as $\eta=0.005$. For the grid-connected power converter, the number of gradient descent steps per control sampling time is set to $1$. For the wind generator, two configurations are compared with $1$ and $10$ gradient descent steps per control sampling time.
V-B Simulations on a grid-connected power converter
The grid-connected power converter system considered in this study is shown in Fig. 1, where we employ the DeePO controller. The objective of the DeePO controller is to regulate the deviations of active and reactive power from their references to zero with minimal control effort, thereby effectively mitigating power oscillations caused by grid disturbances.
Fig. 2 shows the time-domain responses of the power converter. At $t=1.0~{}\mathrm{s}$, we change the short circuit ratio of the system from $5$ to $2.21$ to emulate an event in the power grid, for example, tripping of transmission lines. It can be seen that the converter starts to oscillate after the disturbance, which is caused by the interactions among PLL, current/power control loops, and the weak power grid [2]. Such oscillations are undesirable, as they may trigger resonance in the system, increase the risk of equipment failure, and even result in grid collapse.
Figure 2: Time-domain responses of the grid-connected power converter. The DeePO is activated at $t=3.5\,\mathrm{s}$. with DeePO; without DeePO.
After the oscillation is observed, we inject band-limited white noise signals into the system through the two input channels from $t=2.0\,\mathrm{s}$ to $t=3.5\,\mathrm{s}$ to excite the system and collect data. The DeePO controller is activated at $t=3.5\,\mathrm{s}$, providing real-time control inputs that effectively counteract the oscillation. As depicted in Fig. 2, DeePO effectively damps the oscillation and restores the stable operation of the power converter. In contrast, the responses of the power converter without DeePO are shown as the grey lines in Fig. 2, where the system exhibits sustained active and reactive power oscillations, indicating a low stability margin and increasing the risk of equipment damage and potential cascading failures in the power system.
V-C Simulations on a direct-drive wind generator
We consider now a direct-drive wind generator with a high-fidelity model implemented in MATLAB/Simulink (2023b) [21]. The wind generator model here is more complicated than the converter model in the last subsection, as it not only has a converter as the grid interface, but also considers the generator dynamics on the DC side. The simulation model includes detailed turbine dynamics, flux dynamics, filters, speed control, pitch control, converter control, and maximum power point tracking (MPPT). The grid is modeled as a voltage source behind a transmission line with unknown impedance, which impacts the wind generator in a closed-loop manner. The whole system is illustrated in Fig. 3. Due to proprietary manufacturer models and the complexity of the power grid, its exact dynamical model can hardly be derived or identified. In what follows, we apply the DeePO method in Algorithm 2 to stabilize the wind generator.
Figure 3: The application of DeePO to stabilize a direct-drive wind generator.
Figure 4: The time-domain responses of the wind generator: (a) active power and (b) reactive power. The DeePO is activated at $t=6.0\,\mathrm{s}$. with DeePO (1 gradient descent step per control cycle); with DeePO (10 gradient descent steps per control cycle); without DeePO.
Fig. 4 shows the time-domain responses of the wind generator under different configurations of DeePO. At $t=1.0\,\mathrm{s}$, the short circuit ratio is changed from $2.23$ to $2$, simulating a grid event, such as the tripping of transmission lines. Following the disturbance, the wind generator begins to oscillate.
Then, we inject band-limited white noise to collect input-output trajectory data from $t=3.0\,\mathrm{s}$ to $t=6.0\,\mathrm{s}$. At $t=6.0\,\mathrm{s}$, DeePO is activated and successfully provides real-time control inputs, eliminating the oscillations in both configurations, as shown in Fig. 4. In the configuration with fewer gradient descent steps per control cycle (blue line), the convergence of active and reactive power is slower, with acceptable stabilization performance. By comparison, using $10$ steps (orange line) enables faster damping of oscillations. This illustrates a trade-off between computational efficiency and control performance: increasing the gradient steps improves convergence but requires higher computational effort. By comparison, the wind generator system exhibits a low stability margin without DeePO (grey line in Fig. 4), with persistent oscillations in active and reactive power after $t=6.0\,\mathrm{s}$. These persistent oscillations endanger the wind generator itself (as mechanical resonance may be triggered) and also the secure operation of power systems. We observe that the damping performance of DeePO is similar to the performance of DeePC (c.f. Fig. 14 in [21]), but the computational burden is significantly less (since DeePC requires solving a quadratic program online).
Figure 5: The time-domain responses of the wind generator: (a) active power and (b) reactive power. with DeePO controller; with non-adaptive controller.
In time-varying environments, non-adaptive controllers that rely solely on historical data may fail to maintain effective control. In contrast, DeePO adapts by continuously learning from real-time data, allowing it to better accommodate time-varying system conditions. Fig. 5 shows the time-domain responses of the wind generator using a non-adaptive controller (a fixed controller that achieves similar performance to DeePO before $t=9.0\,\mathrm{s}$) and the DeePO controller, respectively, under parameter changes and disturbances. Before $t=9.0\,\mathrm{s}$, both controllers have converged to their respective optimal gains based on previously collected data. At $t=9.0\,\mathrm{s}$, the DC voltage control parameters are intentionally reduced to 0.9 times of their original values, simulating a sudden change in the wind generator system. While this change does not affect the equilibrium point, it modifies the dynamic characteristics of the system. From $t=9.0\,\mathrm{s}$ to $t=21.0\,\mathrm{s}$, the DeePO controller continues to update its control strategy by leveraging real-time data, whereas the non-adaptive controller retains its original control settings without adaptation. To ensure that the collected data are informative to update the DeePO controller, noise injection is increased between $t=10.0\,\mathrm{s}$ and $t=17.0\,\mathrm{s}$ to better excite the system and capture its dynamic characteristics.
At $t=21.0\,\mathrm{s}$, a small disturbance is introduced by injecting a q-axis voltage perturbation into the PLL. As shown in Fig. 5, the DeePO controller quickly damps out oscillations, demonstrating significantly better response performance compared to the non-adaptive controller. This improvement can be attributed to the real-time adaptation of the control policy in DeePO, which enables it to perceive changes in system dynamics and maintain superior performance under disturbances.
VI Conclusion
In this paper, we proposed a data-enabled policy optimization (DeePO) method, which is direct data-driven, adaptive, based on online input-output data, and recursive in terms of implementation. We demonstrated via simulations the effectiveness of DeePO in stabilizing and mitigating undesired oscillations in a grid-connected power converter and a direct-drive wind generator. In particular, we observed that DeePO has satisfactory performance in stabilization, and shows a similar damping ratio to data-enabled predictive control (DeePC). Nevertheless, DeePO is computationally more efficient than DeePC, as it is gradient-based and does not require solving optimization problems online.
Future work includes the extension of DeePO to achieve other control objectives (e.g., reference tracking) and to other types of systems (e.g., slowly time-varying systems).
Appendix A Proof of Theorem 1
The state can be represented with system dynamics and history trajectories as
$x_{t}$ $=A^{n}x_{t-n}+\mathcal{C}{u}_{t,n}+\mathcal{C}_{w}w_{t,n}$ (12)
${y}_{t,n}$ $=\mathcal{O}x_{t-n}+\mathcal{T}{u}_{t,n}+\mathcal{T}_{w}w_{t,n}+v_{t,n}.$
Since $\mathcal{O}$ has full column rank, it has a unique left pseudo inverse $\mathcal{O}^{\dagger}=(\mathcal{O}^{\top}\mathcal{O})^{-1}\mathcal{O}^{\top}$. Then, it follows immediately from (12) that
$x_{t}$ $=(\mathcal{C}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}){u}_{t,n}+A^{n}\mathcal{O}^{\dagger}{y}_{t,n}$
$~{}~{}~{}+(\mathcal{C}_{w}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}_{w})w_{t,n}-A^{n}\mathcal{O}^{\dagger}v_{t,n},$
and $y_{t}=Cx_{t}+v_{t}=Sz_{t}+d_{t}.$ Thus, a non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization).
The proof for controllablity of (LABEL:equ:nonnimimal_realization) follows the same vein of that of the key reachability lemma in [30, Lemma 3.4.7]. Note that their relatively prime condition holds for the system (1) since the length of the past trajectory used to represent the state matches the system order $n$.

### v2

@@
 Abstract
-Power electronic converters are gradually becoming the main components of modern power systems due to the increasing integration of renewable energy sources.
+Power electronic converters are becoming the main components of modern power systems due to the increasing integration of renewable energy sources.
 However, power converters may become unstable when interacting with the complex and time-varying power grid.
-To deal with this problem, an adaptive control design scheme for power converters is preferable, which can capture the closed-loop dynamical interaction between the converter and the grid via online data.
-In this paper, we propose an adaptive data-driven control method, called data-enabled policy optimization (DeePO), to stabilize power converters by using only online input-output data.
+In this paper, we propose an adaptive data-driven control method to stabilize power converters by using only online input-output data.
 Our contributions are threefold.
-First, we propose a covariance parameterization of partially observed linear systems with input-output data.
-Second, we develop a DeePO algorithm, which updates the parameterized policy with data-based gradient descent to achieve computationally efficient adaptive control.
-Third, we use high-fidelity simulations to verify that DeePO can effectively stabilize grid-connected power converters and quickly adapt to the changes in the power grid.
+First, we reformulate the output-feedback control problem as a state-feedback linear quadratic regulator (LQR) problem with a controllable non-minimal state, which can be constructed from past input-output signals.
+Second, we propose a data-enabled policy optimization (DeePO) method for this non-minimal realization to achieve efficient output-feedback adaptive control.
+Third, we use high-fidelity simulations to verify that the output-feedback DeePO can effectively stabilize grid-connected power converters and quickly adapt to the changes in the power grid.
 I Introduction
@@
 We call this method Data-enabled Policy Optimization (DeePO), where the policy is parameterized with sample covariance of input-state data and updated using gradient methods.
-DeePO has a recursive algorithmic implementation and hence is computationally efficient.
-From a theoretical perspective, it meets provable stability and convergence guarantees for linear time-invariant systems.
-All these key features render DeePO an ideal data-driven control method for power converters.
-In this paper, we propose a DeePO method to mitigate oscillations in power converter systems.
-Since the power converter system is only partially observed, i.e., the full state is unmeasurable, we first extend the covariance parameterization of the system with input-state data in [23, 24] to that with input-output data.
-This is achieved by regarding a finite length of past input-output trajectory as the state and using the corresponding covariance to parameterize the closed-loop system.
-Then, we propose the DeePO algorithm for direct adaptive control of partially observed systems.
-Finally, we apply the DeePO algorithm to stabilize power converter and direct-drive wind generator systems, both of which are unknown, state-unmeasurable, and encounter sudden change of dynamics due to grid changes.
-Simulation results show that DeePO enables efficient online adaptation and effectively prevents instabilities in grid-connected power converters.
+DeePO is computationally efficient, meets provable stability and convergence guarantees for linear time-invariant systems, and has successful real-world applications [25].
+However, the DeePO method [24] works only on the state-feedback control problem, which is not the case for the power converter system.
+In this paper, we propose an output-feedback DeePO method to mitigate oscillations in power converter systems.
+To this end, we first reformulate the output-feedback control problem as a state-feedback LQR problem with a controllable non-minimal state, which can be constructed from input-output signals.
+This is achieved by the state reduction method in [26].
+Then, we apply DeePO to this controllable non-minimal realization to achieve output-feedback adaptive control.
+Finally, we apply the output-feedback DeePO algorithm to stabilize power converter and direct-drive wind generator systems, both of which are unknown, state-unmeasurable, and encounter sudden change of dynamics due to grid changes.
+Simulation results show that output-feedback DeePO enables efficient online adaptation and effectively prevents instabilities in grid-connected power converters.
 Compared with our previous works applying data-enabled predictive control (DeePC) to power converters [21, 27], DeePO offers a significantly reduced online computational burden, making it more suitable for scenarios where the processor cannot solve a quadratic program in real time.
@@
 Section II formulates the stabilization problem of power converter systems.
-Section III introduces the adaptive LQR control and the DeePO method.
-Section IV extends the DeePO algorithm to input-output systems.
+Section III proposes a controllable non-minimal realization of the output feedback system.
+Section IV proposes an output-feedback DeePO algorithm for adaptive control.
 Section V performs simulations on the converter systems.
-Conclusion is made in Section VI.
+Conclusions are made in Section VI.
 Notation.
@@
 Without loss of generality, consider the state-space model linearized at the origin
-$x_{t+1}$ $=Ax_{t}+Bu_{t}+w_{t},$ (1)
+$x_{t+1}$ $=Ax_{t}+Bu_{t}+d_{t},$ (1)
 $y_{t}$ $=Cx_{t}+v_{t},$
 where $x_{t}\in\mathbb{R}^{n}$ is the state variable (could be unmeasurable, e.g., variables in the power grid side), $u_{t}\in\mathbb{R}^{m}$ is the control input (e.g., additional signals added to the current references; see Fig.
-1), $y_{t}\in\mathbb{R}^{d}$ is the output (e.g., active and reactive power in our setting), $w_{t}$ is the process noise, and $v_{t}$ is the measurement noise.
+1), $y_{t}\in\mathbb{R}^{p}$ is the output (e.g., active and reactive power in our setting), $d_{t}$ is the process noise, and $v_{t}$ is the measurement noise.
 The unknown system $(A,B,C)$ is controllable and observable, but it may be subject to sudden impedance changes due to external grid changes, e.g., changes of $Z_{\rm grid}$ in Fig.
@@
   • unknown model: Power converter systems often exhibit high-order characteristics, rendering it difficult to establish exact dynamical models.
-Specifically, the exact dynamics of the DC source and the converter are usually hard to obtain (and therefore unknown) due to proprietary manufacturer models and their complexity, e.g., the complex mechanical behaviors of wind turbines.
-Moreover, the power grid is also unknown when designing a controller for the converter, as the grid is large-scale and time-varying (therefore hard to model).
+Moreover, the power grid is also unknown when designing a controller for the converter, as the grid is large-scale and time-varying.
   • measurement limitations: Internal states are not measurable and hence we can only use input-output information for the control design.
@@
 It is data-driven and does not involve any explicit SysID.
-Moreover, DeePO uses online closed-loop data to adaptively update the control policy, making it an ideal method to deal with time-varying dynamics.
-In the sequel, we first recapitulate the DeePO method with input-state data in [23, 24].
-Then, we extend DeePO to the input-output setting.
-Finally, we discuss its implementation on power converter systems and perform simulations to validate the effectiveness of DeePO.
-III Direct adaptive LQR control via data-enabled policy optimization
-In this section, we first introduce the linear quadratic regulator (LQR). Then, we recapitulate data-enabled policy optimization for adaptive learning of the LQR [23, 24].
-III-A The linear quadratic regulator
-Consider a linear time-invariant system with fully measurable state
+Moreover, DeePO uses online closed-loop data to adaptively update the control policy, making it suitable to deal with time-varying dynamics.
+However, the DeePO method [24] works only on the state-feedback setting, which is not the case for the system (1).
+Leveraging the results from [26], we first propose a controllable non-minimal realization of (1), whose state is measurable from past input-output signals.
+Then, we propose an output-feedback DeePO algorithm for this non-minimal realization.
+Finally, we discuss its implementation on power converter systems and perform simulations to validate the effectiveness of the output-feedback DeePO algorithm.
+III A controllable non-minimal realization of the output-feedback system
+In this section, we leverage the results from [26] to find a controllable non-minimal realization of (1), where the state can be measured from input-output signals.
+III-A A non-minimal controllable state
+We first consider the system (1) without noise, i.e., $d_{t}=0,v_{t}=0$.
+We assume that the system order $n$ and lag $l$ is unknown, but we have prior knowledge on an upper bound of the lag $\bar{l}\geq l$.
+Since the state is unmeasurable, we represent (1) with a non-minimal realization using input and output signals.
+Denote the input trajectory, the output trajectory, and their stack from time $t-\bar{l}$ to $t-1$ as
+${u}_{t,\bar{l}}=\begin{bmatrix}u_{t-1}\\ \vdots\\ u_{t-\bar{l}}\end{bmatrix},~{}{y}_{t,\bar{l}}=\begin{bmatrix}y_{t-1}\\ \vdots\\ y_{t-\bar{l}}\end{bmatrix},~{}{\xi}_{t}=\begin{bmatrix}{u}_{t,\bar{l}}\\ {y}_{t,\bar{l}}\end{bmatrix},$
+respectively. Define the extended observability matrix
+$\mathcal{O}=\begin{bmatrix}CA^{\bar{l}-1}\\ \vdots\\ CA\\ C\end{bmatrix},$
+the controllability matrix ${}\mathcal{C}=[B~{}~{}AB~{}~{}\cdots~{}~{}A^{\bar{l}-1}B]$, and the Toeplitz matrices capturing the impulse response
+$\mathcal{T}=\begin{bmatrix}0&CB&CAB&\cdots&CA^{\bar{l}-2}B\\ 0&0&CB&\cdots&CA^{\bar{l}-3}B\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&CB\\ 0&0&0&0&0\end{bmatrix}.$
+Let $S=[C(\mathcal{C}-A^{\bar{l}}O^{\dagger}\mathcal{T}),CA^{\bar{l}}O^{\dagger}]$. Then, we have the following results.
+Lemma 1 (A non-minimal realization)
+A non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization) shown at the bottom of this page.
+Proof:
+The state can be represented with system dynamics and past input-output trajectories as
+$x_{t}$ $=A^{\bar{l}}x_{t-{\bar{l}}}+\mathcal{C}{u}_{t,{\bar{l}}}$ (2)
+${y}_{t,{\bar{l}}}$ $=\mathcal{O}x_{t-{\bar{l}}}+\mathcal{T}{u}_{t,{\bar{l}}}.$
+Since ${\bar{l}}\geq l$, the extended observability matrix $\mathcal{O}$ has full column rank, and it has a unique left pseudo inverse $\mathcal{O}^{\dagger}=(\mathcal{O}^{\top}\mathcal{O})^{-1}\mathcal{O}^{\top}$.
+Then, it follows immediately from (2) that $x_{t}=(\mathcal{C}-A^{\bar{l}}\mathcal{O}^{\dagger}\mathcal{T}){u}_{t,{\bar{l}}}+A^{\bar{l}}\mathcal{O}^{\dagger}{y}_{t,{\bar{l}}}$ and $y_{t}=Cx_{t}=S\xi_{t}.$ Thus, a non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization).
+However, for multiple-output systems with $p>1$, the non-minimal realization (LABEL:equ:nonnimimal_realization) is generally not controllable. Moreover, the corresponding input-state data matrix
+$\begin{bmatrix}u_{0}&u_{1}&\dots&u_{t-1}\\ \xi_{0}&\xi_{1}&\dots&\xi_{t-1}\end{bmatrix}\in\mathbb{R}^{(m(\bar{l}+1)+p\bar{l})\times t}$ (3)
+will never have full row rank [26], which precludes the application of DeePO on the non-minimal realization (LABEL:equ:nonnimimal_realization) [24].
+In fact, it can be shown that the maximal rank of (3) can only be $m(\bar{l}+1)+n$ [26].
+To ensure the full rank condition of data, we adopt the approach [26] that constructs an reduced non-minimal state for (1). The following lemma is a direct implication of [26, Theorem 4].
+Lemma 2 (A controllable non-minimal realization)
+Let $d_{t}=0,v_{t}=0$ for (1).
+Then, there exists a full row rank permutation matrix $T\in\mathbb{R}^{(m\bar{l}+n)\times(m\bar{l}+p\bar{l})}$ and $(A_{z},B_{z},C_{z})$ such that a controllable non-minimal state-space representation is given by
+$z_{t+1}$ $=A_{z}z_{t}+B_{z}u_{t}$ (4)
+$y_{t}$ $=C_{z}z_{t}$
+with a non-minimal state
+$z_{t}=T\xi_{t}.$ (5)
+Since the realization (4) is controllable, we can choose persistently exciting (PE) inputs of order $\bar{l}+1+n$ such that
+$D_{0,t}=\begin{bmatrix}u_{0}&u_{1}&\dots&u_{t-1}\\ z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{(m(\bar{l}+1)+n)\times t}$ (6)
+has full row rank [26]. Hence, we can design data-driven state-feedback controller for (4), where $z_{t}$ can be computed with the permutation $T$ and input-output signals.
+Next, we show how to obtain the permutation matrix $T$ from a batch of input-output data [26].
+III-B Computing the permutation matrix from input-output data
+When there is no noise in the system (1), the permutation matrix $T$ can be solved directly from a batch of input-output data [26].
+Specifically, suppose that we have the data generated by PE input of order $\bar{l}+1+n$
+$\Xi_{0,t}=\begin{bmatrix}{\xi}_{0}&{\xi}_{1}&\dots&{\xi}_{t-1}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+p\bar{l})\times t}.$ (7)
+Then, $T\in\mathbb{R}^{(m\bar{l}+n)\times(m\bar{l}+p\bar{l})}$ is the permutation matrix such that $T\Xi_{0,t}$ has full row rank $m\bar{l}+n$.
+When $n$ is unknown, we can easily separate linearly independent rows by, e.g., Gaussian elimination, and further obtain $T$.
+Moreover, the system order $n$ can be obtained as the number of rows in $T$.
+When the noise $d_{t}$ and $v_{t}$ in system (1) are not zero, the input-state data matrix (3) of system (LABEL:equ:nonnimimal_realization) can be full row rank.
+In this case, we use singular value decomposition (SVD) for $\Xi_{0,t}$
+$\Xi_{0,t}=\begin{bmatrix}U_{r}&U_{l}\end{bmatrix}\begin{bmatrix}\Lambda_{r}&0\\ 0&\Lambda_{l}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\ V_{l}^{\top}\end{bmatrix},$
+where $\Lambda_{r}$ is the singular value matrix with largest $r$ singular values.
+The number of rows $r$ of the permutation matrix $T$ should be $m\bar{l}+n$, which is consistent with Lemma 2.
+Since $n$ is unknown, we can select $r>m\bar{l}$ such that there is a clear distinction between the largest $r$ singular values and the remaining ones (corresponding to noise).
+Then, the permutation matrix $T$ is given by
+$T=\Lambda_{r}^{-1}U_{r}^{\top},$
+i.e., it selects the orthogonal basis for the row space of $\Xi_{0,t}$ corresponding to largest $r$ singular values, i.e., $V_{r}^{\top}=T\Xi_{0,t}.$
+Once we obtain the permutation matrix $T$, we can measure the non-minimal state $z_{t}$ from input-output signals via (5).
+Moreover, the input-state data matrix (6) has full rank, which enables us to use DeePO for the realization (4) for adaptive control.
+IV Output-feedback data-enabled policy optimization for direct adaptive control
+In this section, we first introduce the data-driven LQR formulation with covariance parameterization for the non-minimal realization (4).
+Based on this, we propose a DeePO method for direct adaptive control of (1) using input-output data.
+IV-A Direct data-driven LQR design of (4) with covariance parameterization
+Consider the controllable non-minimal realization (4) with noise
 […]
-Here, $h_{t}$ is the performance signal of interest, $(A,B)$ is controllable, and the weighting matrices $(Q,R)$ are positive definite.
-The LQR problem is phrased as finding a state-feedback gain $K\in\mathbb{R}^{m\times n}$ that minimizes the $\mathcal{H}_{2}$-norm of the transfer function $\mathscr{T}(K):w\rightarrow h$ of the closed-loop system
-$\begin{bmatrix}x_{t+1}\\ h_{t}\end{bmatrix}=\begin{bmatrix}A+BK&I_{n}\\ \hline\cr\begin{bmatrix}Q^{1/2}\\ R^{1/2}K\end{bmatrix}&0\end{bmatrix}\begin{bmatrix}x_{t}\\ w_{t}\end{bmatrix}.$
-When $A+BK$ is stable, it holds that [26]
-$\|\mathscr{T}(K)\|_{2}^{2}=\text{Tr}((Q+K^{\top}RK)\Sigma_{K})=:C(K),$ (3)
+Here, $h_{t}$ is the performance signal of interest and the weighting matrices $(Q,R)$ are positive definite.
+The LQR problem is phrased as finding a state-feedback gain $K\in\mathbb{R}^{m\times(m\bar{l}+n)}$ that minimizes the $\mathcal{H}_{2}$-norm of the transfer function $\mathscr{T}(K):w\rightarrow h$ of the closed-loop system
+$\begin{bmatrix}z_{t+1}\\ h_{t}\end{bmatrix}=\begin{bmatrix}A_{z}+B_{z}K&I_{n}\\ \hline\cr\begin{bmatrix}Q^{1/2}\\ R^{1/2}K\end{bmatrix}&0\end{bmatrix}\begin{bmatrix}z_{t}\\ d_{t}\end{bmatrix}.$
+When $A_{z}+B_{z}K$ is stable, it holds that [28]
+$\|\mathscr{T}(K)\|_{2}^{2}=\text{Tr}((Q+K^{\top}RK)\Sigma_{K})=:J(K),$ (9)
 where $\Sigma_{K}$ is the closed-loop state covariance matrix obtained as the positive definite solution to the Lyapunov equation
-$\Sigma_{K}=I_{n}+(A+BK)\Sigma_{K}(A+BK)^{\top}.$ (4)
-We refer to $C(K)$ as the LQR cost and to (3)-(4) as a policy parameterization of the LQR.
-It is well-known that the optimal LQR gain $K^{*}$ is unique and can be found by, e.g., solving an algebraic Riccati equation with $(A,B)$ [26].
-When $(A,B)$ is unknown, data-driven methods learn the LQR gain from input-state data.
-In the sequel, we recall the DeePO approach proposed in our previous work [23, 24].
-III-B DeePO for direct adaptive LQR control
-Policy optimization (PO) refers to a class of direct design methods, where the policy is parameterized and recursively updated using gradient methods [27].
-In particular, DeePO uses the sample covariance of input-state data to parameterize the policy [24].
-Consider the $t$-long time series of states, inputs, noises, and successor states
-$X_{0,t}$ $:=\begin{bmatrix}x_{0}&x_{1}&\dots&x_{t-1}\end{bmatrix}\in\mathbb{R}^{n\times t},$
+$\Sigma_{K}=I_{n}+(A_{z}+B_{z}K)\Sigma_{K}(A_{z}+B_{z}K)^{\top}.$ (10)
+We refer to $J(K)$ as the LQR cost and to (9)-(10) as a policy parameterization of the LQR.
+When the model parameters are known, the optimal LQR gain $K^{*}$ is unique and can be found by, e.g., solving an algebraic Riccati equation [28].
+Since $(A_{z},B_{z})$ is unknown, data-driven methods learn the LQR gain from input-state data. Consider the $t$-long time series of states, inputs, noises, and successor states of (4)
+$Z_{0,t}$ $:=\begin{bmatrix}z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+n)\times t},$
 $U_{0,t}$ $:=\begin{bmatrix}u_{0}&u_{1}&\dots&u_{t-1}\end{bmatrix}\in\mathbb{R}^{m\times t},$
-$W_{0,t}$ $:=\begin{bmatrix}w_{0}&w_{1}&\dots&w_{t-1}\end{bmatrix}\in\mathbb{R}^{n\times t},$
-$X_{1,t}$ $:=\begin{bmatrix}x_{1}&x_{2}&\dots&x_{t}\end{bmatrix}\in\mathbb{R}^{n\times t},$
+$W_{0,t}$ $:=\begin{bmatrix}w_{0}&w_{1}&\dots&w_{t-1}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+n)\times t},$
+$Z_{1,t}$ $:=\begin{bmatrix}z_{1}&z_{2}&\dots&z_{t}\end{bmatrix}\in\mathbb{R}^{(m\bar{l}+n)\times t},$
 which satisfy the system dynamics
-$X_{1,t}=AX_{0,t}+BU_{0,t}+W_{0,t}.$ (5)
-Assuming that the data is persistently exciting (PE) [28], i.e., the block matrix of input and state data $D_{0,t}:=[U_{0,t}^{\top},X_{0,t}^{\top}]^{\top}$ has full row rank
-$\text{rank}(D_{0,t})=m+n,$ (6)
-the covariance parameterization[24] of $K$ is given by
-$\begin{bmatrix}K\\ I_{n}\end{bmatrix}=\frac{1}{t}D_{0,t}D_{0,t}^{\top}V=:\Phi_{t}V,$ (7)
-where $\Phi_{t}\succ 0$ denotes the sample covariance of input-state data, and $V\in\mathbb{R}^{(n+m)\times n}$ is the parameterized policy.
-With the covariance parameterization (7), the LQR problem (3)-(4) can be expressed by raw data matrices $(X_{0,t},U_{0,t},X_{1,t})$ and the optimization matrix $V$.
-For brevity, let $\overline{X}_{0,t}=X_{0,t}D_{0,t}^{\top}/t$ and $\overline{U}_{0,t}=U_{0,t}D_{0,t}^{\top}/t$ be a partition of $\Phi_{t}$, and let $\overline{W}_{0,t}=W_{0,t}D_{0,t}^{\top}/t$ be the noise-state-input covariance, and finally define the covariance with respect to the successor state $\overline{X}_{1,t}=X_{1,t}D_{0,t}^{\top}/t$, Note that the dimension of the covariance matrices does not depend on $t$.
+$Z_{1,t}=A_{z}Z_{0,t}+B_{z}U_{0,t}+W_{0,t}.$ (11)
+Assume that the input data is PE of order $\bar{l}+1+n$. Then, the block matrix $[U_{0,t}^{\top},Z_{0,t}^{\top}]^{\top}$ has full row rank. Define the sample covariance of input-state data as
+$\Phi_{t}:=\frac{1}{t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top},$ (12)
+which is positive definite due to the full rank condition. Then, we can use sample covariance to parameterize the policy
+$\begin{bmatrix}K\\ I_{(m\bar{l}+n)}\end{bmatrix}=\Phi_{t}V,$ (13)
+where $V\in\mathbb{R}^{(m\bar{l}+m+n)\times(m\bar{l}+m+n)}$.
+With the covariance parameterization (13), the LQR problem (9)-(10) can be expressed by raw data matrices $(Z_{0,t},U_{0,t},Z_{1,t})$ and the optimization matrix $V$.
+For brevity, let $\overline{Z}_{0,t}=Z_{0,t}D_{0,t}^{\top}/t$ and $\overline{U}_{0,t}=U_{0,t}D_{0,t}^{\top}/t$ be a partition of $\Phi_{t}$, and let $\overline{W}_{0,t}=W_{0,t}D_{0,t}^{\top}/t$ be the noise-state-input covariance, and finally define the covariance with respect to the successor state $\overline{Z}_{1,t}=Z_{1,t}D_{0,t}^{\top}/t$.
 Then, the closed-loop matrix can be written as
 […]
-Following the certainty-equivalence principle [14], we disregard the unmeasurable $\overline{W}_{0,t}$ and use $\overline{X}_{1,t}V$ as the closed-loop matrix.
-After substituting $A+BK$ with $\overline{X}_{1,t}V$ in (3)-(4) and leveraging (7), the LQR problem becomes
-$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V)\Sigma_{t}(V)\right),$ (8)
-subject $\text{subject to}~{}~{}\overline{X}_{0,t}V=I_{n},$
-where $\Sigma_{t}(V)=I_{n}+\overline{X}_{1,t}V\Sigma_{t}(V)V^{\top}\overline{X}_{1,t}^{\top}$ is a covariance parameterization of (4), and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$.
+Following the certainty-equivalence principle [14], we disregard the unmeasurable $\overline{W}_{0,t}$ and use $\overline{Z}_{1,t}V$ as the closed-loop matrix.
+After substituting $A+BK$ with $\overline{Z}_{1,t}V$ in (9)-(10) and leveraging (13), the LQR problem becomes
+$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V)\Sigma_{t}\right),$ (14)
+subject $\text{subject to}~{}~{}\overline{Z}_{0,t}V=I_{m\bar{l}+n},$
+where $\Sigma_{t}=I_{m\bar{l}+n}+\overline{Z}_{1,t}V\Sigma_{t}V^{\top}\overline{Z}_{1,t}^{\top}$ is a covariance parameterization of (10), and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$.
 We refer to (14) as the direct data-driven LQR problem, which does not involve any explicit SysID.
-The DeePO algorithm uses online gradient descent of (8) to recursively update $V$.
-The details are presented in Algorithm 1, which alternates between control (line 2) and policy update (line 3-5).
-The initial stabilizing policy can be either user-specified or computed by solving (8) with offline data.
-At time $t$, we apply the linear state feedback policy $u_{t}=K_{t}x_{t}+e_{t}$ for control, where $e_{t}$ is a probing noise used to ensure the PE rank condition (6).
-We use online projected gradient descent in (10) to update the parameterized policy, where the projection $\Pi_{\overline{X}_{0,t+1}}:=I_{n+m}-\overline{X}_{0,t+1}^{\dagger}\overline{X}_{0,t+1}$ onto the nullspace of $\overline{X}_{0,t+1}$ is to ensure the subspace constraint in (8).
+IV-B Output-feedback DeePO
+Policy optimization (PO) refers to a class of direct design methods, where the policy is parameterized and recursively updated using gradient methods [29].
+In particular, the DeePO algorithm uses online gradient descent of (14) to recursively update $V$.
+The details are presented in Algorithm 1.
+Given offline data, we first compute the permutation matrix $T$.
+At time $t$, we apply the linear state feedback policy $u_{t}=K_{t}z_{t}+e_{t}$ for control, where $e_{t}\in\mathbb{R}^{m}$ is a probing noise used to ensure the PE rank condition.
+We use online projected gradient descent to update the parameterized policy, where the projection $\Pi_{\overline{Z}_{0,t+1}}:=I_{m(\bar{l}+1)+n}-\overline{Z}_{0,t+1}^{\dagger}\overline{Z}_{0,t+1}$ onto the nullspace of $\overline{Z}_{0,t+1}$ is to ensure the subspace constraint in (14).
+Define the feasible set of (14) (i.e., the set with stable closed-loop matrices) as $\mathcal{S}_{t}:=\{V\mid\overline{Z}_{0,t}V=I_{m\bar{l}+n},\rho(\overline{Z}_{1,t}V)<1\}$.
 The gradient can be computed by the following lemma.
-Define the feasible set of (8) (i.e., the set with stable closed-loop matrices) as $\mathcal{S}_{t}:=\{V\mid\overline{X}_{0,t}V=I_{n},\rho(\overline{X}_{1,t}V)<1\}$.
-Lemma 1 ([24])
+Lemma 3 ([24, Lemma 2])
 For $V\in\mathcal{S}_{t}$, the gradient of $J_{t}(V)$ with respect to $V$ is given by
-$\nabla J_{t}(V)=2\left(\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}+\overline{X}_{1,t}^{\top}P_{t}\overline{X}_{1,t}\right)V\Sigma_{t}(V),$ (9)
+$\nabla J_{t}(V)=2\left(\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}+\overline{Z}_{1,t}^{\top}P_{t}\overline{Z}_{1,t}\right)V\Sigma_{t},$ (15)
 where $P_{t}$ satisfies the Lyapunov equation
-$P_{t}=Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V+V^{\top}\overline{X}_{1,t}^{\top}P_{t}\overline{X}_{1,t}V.$
-DeePO is direct and adaptive in the sense that it directly uses online closed-loop data to update the policy.
+$P_{t}=Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V+V^{\top}\overline{Z}_{1,t}^{\top}P_{t}\overline{Z}_{1,t}V.$
+By Lemma 3, the gradient can be computed by solving two Lyapunov equations.
+The stepsize $\eta_{t}$ should be chosen based on the signal-to-noise ratio (SNR) of the online data.
+Specifically, when the SNR is high, the gradient direction is more reliable, allowing for a larger stepsize.
+Conversely, when the SNR is low, a smaller stepsize is necessary to prevent the policy from deviating from the stability region.
+Accordingly, we set the stepsize as
+$\eta_{t}=\frac{\eta_{0}}{\left\|\overline{U}_{0,t}\Pi_{\overline{Z}_{0,t}}\overline{U}_{0,t}^{\top}\right\|},~{}t\geq t_{0},$ (16)
+where $\eta_{0}$ is a constant, and the denominator is used to quantify the SNR.
+The DeePO algorithm is direct and adaptive in the sense that it directly uses online closed-loop data to update the policy.
 Thus, it can rapidly adapt to dynamic changes reflected in the data.
@@
 $V_{t+1}=\frac{t+1}{t}\left(V_{t}^{\prime}-\frac{\Phi_{t}^{-1}\phi_{t}\phi_{t}^{\top}V_{t}^{\prime}}{t+\phi_{t}^{\top}\Phi_{t}^{-1}\phi_{t}}\right),$
-where $\phi_{t}=[u_{t}^{\top},x_{t}^{\top}]^{\top}$, and $\Phi_{t}^{-1}$ and $V_{t}^{\prime}$ are given from the last iteration.
+where $\phi_{t}=[u_{t}^{\top},z_{t}^{\top}]^{\top}$, and $\Phi_{t}^{-1}$ and $V_{t}^{\prime}$ are given from the last iteration.
 Theoretically, it is shown that under mild assumptions the policy $\{K_{t}\}$ converges to the optimal LQR gain.
 We refer to [24, Section IV] for detailed discussions.
-The DeePO framework developed so far considers the state feedback setting.
-However, the state of power converter system (1) is not measurable, and only input-output data is available for control design.
-Such an system is typically referred to as partially observed [29].
-Next, we extend DeePO to partially observed systems for adaptive control using input-output data.
-Algorithm 1 DeePO for direct adaptive LQR control
-1:Offline data $(X_{0,t_{0}},U_{0,t_{0}},X_{1,t_{0}})$, an initial policy $K_{t_{0}}$, and a stepsize $\eta$.
-2:for $t=t_{0},t_{0}+1,\dots$ do
-3:     Apply $u_{t}=K_{t}x_{t}+e_{t}$ and observe $x_{t+1}$.
-4:     Update covariance matrices $\Phi_{t+1}$ and $\overline{X}_{1,t+1}$.
-5:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
+Algorithm 1 Output-feedback DeePO
+1:Offline data $(u_{-\bar{l}},y_{-\bar{l}},\dots,u_{t_{0}-1},y_{t_{0}-1},y_{t_{0}})$ and a stepsize $\eta_{t}$.
+2:Computation of the permutation matrix: Constitute the data matrix $\Xi_{0,t_{0}}$ in (7) and perform SVD
+$\Xi_{0,t_{0}}=\begin{bmatrix}U_{r}&U_{l}\end{bmatrix}\begin{bmatrix}\Lambda_{r}&0\\ 0&\Lambda_{l}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\ V_{l}^{\top}\end{bmatrix},$
+where $r>m\bar{l}$ is selected such that there is a clear distinction between the largest $r$ singular values and the remaining ones. Let $T=\Lambda_{r}^{-1}U_{r}^{\top}.$
+3:Compute the initial policy $K_{t_{0}}$ from (14) with offline data.
+4:for $t=t_{0},t_{0}+1,\dots$ do
+5:     Compute $z_{t}=T\xi_{t}$, apply $u_{t}=K_{t}z_{t}+e_{t}$ to the system, and observe $z_{t+1}$.
+6:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
 $V_{t+1}=\Phi_{t+1}^{-1}\begin{bmatrix}K_{t}\\ I_{n}\end{bmatrix}.$
-6:     Update of the parameterized policy: perform one-step projected gradient descent
-$V_{t+1}^{\prime}=V_{t+1}-\eta\Pi_{\overline{X}_{0,t+1}}\nabla J_{t+1}(V_{t+1}),$ (10)
-where the gradient $\nabla J_{t+1}(V_{t+1})$ is given by Lemma 1.
-7:     Gain update: update the control gain by
+7:     Update of the parameterized policy: perform one-step projected gradient descent
+$V_{t+1}^{\prime}=V_{t+1}-\eta_{t}\Pi_{\overline{Z}_{0,t+1}}\nabla J_{t+1}(V_{t+1}).$
+8:     Gain update: update the control gain by
 $K_{t+1}=\overline{U}_{0,t+1}V_{t+1}^{\prime}.$
-8:end for
-IV Output-feedback DeePO for partially observed systems
-In this section, we extend DeePO for direct adaptive control of partially observed systems using input-output data.
-Consider the system (1).
-Since the state is unmeasurable, we represent (1) with a non-minimal realization using input and output signals.
-Denote the input trajectory, the output trajectory, and their stack from time $t-n$ to $t-1$ as
-${u}_{t,n}=\begin{bmatrix}u_{t-1}\\ \vdots\\ u_{t-n}\end{bmatrix},~{}{y}_{t,n}=\begin{bmatrix}y_{t-1}\\ \vdots\\ y_{t-n}\end{bmatrix},~{}{z}_{t}=\begin{bmatrix}{u}_{t,n}\\ {y}_{t,n}\end{bmatrix}$
-respectively. Define the observability matrix
-$\mathcal{O}=\begin{bmatrix}CA^{n-1}\\ \vdots\\ CA\\ C\end{bmatrix},$
-the controllability matrices ${}\mathcal{C}=[B~{}~{}AB~{}~{}\cdots~{}~{}A^{n-1}B]$, $\mathcal{C}_{w}=[I~{}~{}A~{}~{}\cdots~{}~{}A^{n-1}],$ and the Toeplitz matrices capturing the impulse and disturbance response
-$\mathcal{T}=\begin{bmatrix}0&CB&CAB&\cdots&CA^{n-2}B\\ 0&0&CB&\cdots&CA^{n-3}B\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&CB\\ 0&0&0&0&0\end{bmatrix},$
-$\mathcal{T}_{w}=\begin{bmatrix}0&C&CA&\cdots&CA^{n-2}\\ 0&0&C&\cdots&CA^{n-3}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&C\\ 0&0&0&0&0\end{bmatrix}.$
-Let $S=[C(\mathcal{C}-A^{n}O^{\dagger}\mathcal{T}),CA^{n}O^{\dagger}]$, $d_{t}=C(\mathcal{C}_{w}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}_{w})w_{t,n}-CA^{n}\mathcal{O}^{\dagger}v_{t,n}+v_{t}$.
-Then, we have the following results, the proof of which is provided in the appendix.
-Theorem 1 (Non-minimal controllable realization)
-A non-minimal controllable realization of (1) is given by (LABEL:equ:nonnimimal_realization) shown at the bottom of the next page.
-Remark 1
-The length of the past trajectory used to represent the state does not necessarily match the system order $n$.
-In fact, it only needs to exceed the system lag (i.e., the observability index), under which the dimension of $z_{t}$ can be reduced.
-Then, we can apply DeePO for the system (LABEL:equ:nonnimimal_realization), whose state is measurable. Define the matrix of input-output data
-$Z_{0,t}=\begin{bmatrix}z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{n(m+d)\times t}.$
-Assume that the data is PE, i.e., the block matrix $[U_{0,t}^{\top},Z_{0,t}^{\top}]^{\top}$ has full row rank.
-Then, the covariance parameterization of $K$ for the system (LABEL:equ:nonnimimal_realization) is given by
-$\begin{bmatrix}K\\ I_{n(m+d)}\end{bmatrix}=\frac{1}{t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top}V=:\Psi_{t}V,$
-where $V\in\mathbb{R}^{n(2m+d)\times n(m+d)}$.
-Define the sample covariance matrices
-$\overline{Z}_{0,t}=\frac{1}{t}Z_{0,t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top},~{}\overline{Z}_{1,t}=\frac{1}{t}Z_{1,t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top}.$
-Denote the realization (LABEL:equ:nonnimimal_realization) as $(A_{z},B_{z},C_{z})$. Following the certainty-equivalence principle and disregarding $d_{t}$, we approximate the closed-loop matrix by
-$A_{z}+B_{z}K\approx\overline{Z}_{1,t}V.$
-Then, the LQR problem of (LABEL:equ:nonnimimal_realization) is parameterized as
-$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q_{1}+V^{\top}\overline{U}_{0,t}^{\top}R_{z}\overline{U}_{0,t}V)\Sigma_{t}(V)\right),$ (11)
-subject $\text{subject to}~{}~{}\overline{Z}_{0,t}V=I_{n},$
-where $Q_{z},R_{z}$ are positive definite matrices with proper dimensions, $\Sigma_{t}(V)=I_{n}+\overline{Z}_{1,t}V\Sigma_{t}(V)V^{\top}\overline{Z}_{1,t}^{\top}$, and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$.
-The DeePO method for partially observed systems is provided in Algorithm 2. Next, we apply Algorithm 2 to stabilize power converter and renewable energy systems.
-Algorithm 2 DeePO for partially observed systems
-1:Offline data $(Z_{0,t_{0}},U_{0,t_{0}},Z_{1,t_{0}})$, an initial policy $K_{t_{0}}$, and a stepsize $\eta$.
-2:for $t=t_{0},t_{0}+1,\dots$ do
-3:     Apply $u_{t}=K_{t}z_{t}+e_{t}$ and observe $z_{t+1}$.
-4:     Update covariance matrices $\Phi_{t+1}$ and $\overline{Z}_{1,t+1}$.
-5:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
-$V_{t+1}=\Psi_{t+1}^{-1}\begin{bmatrix}K_{t}\\ I_{n}\end{bmatrix}.$
-6:     Update of the parameterized policy: perform one-step projected gradient descent
-$V_{t+1}^{\prime}=V_{t+1}-\eta\Pi_{\overline{Z}_{0,t+1}}\nabla J_{t+1}(V_{t+1}).$
-7:     Gain update: update the control gain by
-$K_{t+1}=\overline{U}_{0,t+1}V_{t+1}^{\prime}.$
-8:end for
+9:end for
 V DeePO for stabilization of power converters and renewable energy systems
-In this section, we first discuss the implementation of DeePO on the power converter systems. Then, we perform simulations for a grid-connected power converter and a direct-drive wind generator.
-V-A Implementation of DeePO for power converter systems
+In this section, we first discuss the implementation of output-feedback DeePO on the power converter systems.
+Then, we perform simulations for a grid-connected power converter and a direct-drive wind generator.
+V-A Implementation of output-feedback DeePO for power converter systems
 We consider two typical power converter systems: a grid-connected power converter (with lithium batteries as the DC source) in Fig.
@@
 3, which also uses a converter as the grid interface.
+The maximum lags for the two systems are $\bar{l}=2$ and $4$, respectively.
 Both systems share similar control challenges, including unknown models, measurement limitations, and potential change of dynamics.
@@
 Moreover, since Algorithm 1 has a recursive implementation, the online update of the policy is computationally efficient.
-For both systems, we select the deviations of active and reactive power from their reference values as the outputs of the unknown system, denoted by $y=[P_{E}-P_{0},Q_{E}-Q_{0}]^{\top}$.
-The DeePO method is configured to provide two control inputs that are added to the current references of the system.
+For both systems, the outputs of the unknown dynamics are defined as the deviations of active and reactive power from their reference values: $y=[P_{E}-P_{0},Q_{E}-Q_{0}]^{\top}$.
+The DeePO method is configured to provide two control inputs added to the current references of the system.
 All input and output signals are expressed in per-unit ($\mathrm{p.u.}$) values.
-Since directly measuring internal state variables is challenging, we construct new state variables using input-output information as $z_{t}=[u_{t,T_{\text{ini}}}^{\top},y_{t,T_{\text{ini}}}^{\top}]^{\top}$, where the length of the initial trajectory for constructing new states is set to $T_{\text{ini}}=6$ for both systems.
+Since internal states are not directly measurable, the input-output stack is constructed as $\xi_{t}=[u_{t,\bar{l}}^{\top},y_{t,\bar{l}}^{\top}]^{\top}$.
+The matrix $\Xi_{0,t}$ is built from past trajectories, followed by SVD to determine the reduced dimension $r$.
+A projection matrix $T$ is then computed to obtain the reduced-order state $z_{t}$ from $\xi_{t}$.
 The parameters for the two systems are set as follows:
-  • sampling frequency: The sampling frequency is chosen as $200\,\mathrm{Hz}$.
-  • trajectory length $T$: The trajectory length is set to $300$ for the power converter and $600$ for the wind generator to satisfy the persistency of excitation condition.
-  • weighting matrices: The input weighting matrix is set as $R=I_{m}$, while the output weighting matrix is set as $Q=\text{diag}(I_{mT_{\text{ini}}},100I_{pT_{\text{ini}}})$.
-  • gradient descent parameters: The stepsize for gradient descent is set as $\eta=0.005$.
-For the grid-connected power converter, the number of gradient descent steps per control sampling time is set to $1$.
-For the wind generator, two configurations are compared with $1$ and $10$ gradient descent steps per control sampling time.
+  • The sampling frequency is chosen as $200\,\mathrm{Hz}$.
+  • The trajectory length is set to $300$ for the power converter and $600$ for the wind generator.
+  • The input and state weighting matrices are set as $R=I_{m}$ and $Q=I_{r}$, respectively.
+  • The stepsize for gradient descent is set according to (16) with $\eta_{0}=0.0001$.
 V-B Simulations on a grid-connected power converter
@@
 After the oscillation is observed, we inject band-limited white noise signals into the system through the two input channels from $t=2.0\,\mathrm{s}$ to $t=3.5\,\mathrm{s}$ to excite the system and collect data.
+We then construct the Hankel matrix $\Xi_{0,t}$ using the input-output data from this interval and perform SVD.
+The resulting singular values, in descending order, are 7.075, 2.596, 0.556, 0.496, 0.476, 0.460, 0.249, 0.1867.0752.5960.5560.4960.4760.4600.2490.1867.075,\ 2.596,\ 0.556,\ 0.496,\ 0.476,\ 0.460,\ 0.249,\ 0.1867.075 , 2.596 , 0.556 , 0.496 , 0.476 , 0.460 , 0.249 , 0.186.
+The minimum embedding dimension is $m\bar{l}=4$.
+Since the first six singular values exhibit a clear distinction from the remaining two, we select $r=6$.
+The projection matrix $T$ is then computed based on the leading components, and the reduced-order state is obtained accordingly.
+This representation is used to solve an SDP problem of (14), yielding $K_{t_{0}}$ as the initial policy for DeePO.
 The DeePO controller is activated at $t=3.5\,\mathrm{s}$, providing real-time control inputs that effectively counteract the oscillation.
@@
 In contrast, the responses of the power converter without DeePO are shown as the grey lines in Fig.
-2, where the system exhibits sustained active and reactive power oscillations, indicating a low stability margin and increasing the risk of equipment damage and potential cascading failures in the power system.
+2, where the system exhibits sustained oscillations, indicating a low stability margin and increasing the risk of equipment damage and potential cascading failures in the power system.
 V-C Simulations on a direct-drive wind generator
@@
 Figure 3: The application of DeePO to stabilize a direct-drive wind generator.
-Figure 4: The time-domain responses of the wind generator: (a) active power and (b) reactive power.
-The DeePO is activated at $t=6.0\,\mathrm{s}$.
-with DeePO (1 gradient descent step per control cycle); with DeePO (10 gradient descent steps per control cycle); without DeePO.
+Figure 4: The time-domain responses of the wind generator: (a) active power and (b) reactive power. The DeePO is activated at $t=6.0\,\mathrm{s}$. with DeePO; without DeePO.
 Fig.
@@
 Then, we inject band-limited white noise to collect input-output trajectory data from $t=3.0\,\mathrm{s}$ to $t=6.0\,\mathrm{s}$.
-At $t=6.0\,\mathrm{s}$, DeePO is activated and successfully provides real-time control inputs, eliminating the oscillations in both configurations, as shown in Fig.
-4.
-In the configuration with fewer gradient descent steps per control cycle (blue line), the convergence of active and reactive power is slower, with acceptable stabilization performance.
-By comparison, using $10$ steps (orange line) enables faster damping of oscillations.
-This illustrates a trade-off between computational efficiency and control performance: increasing the gradient steps improves convergence but requires higher computational effort.
-By comparison, the wind generator system exhibits a low stability margin without DeePO (grey line in Fig.
-4), with persistent oscillations in active and reactive power after $t=6.0\,\mathrm{s}$.
-These persistent oscillations endanger the wind generator itself (as mechanical resonance may be triggered) and also the secure operation of power systems.
-We observe that the damping performance of DeePO is similar to the performance of DeePC (c.f.
+Following the same procedure as in the last subsection, we determine the embedding dimension as $r=12$ for this case.
+Based on the data collected in this interval, we solve an SDP of (14) to obtain the initial DeePO policy $K_{t_{0}}$.
+As shown in Fig.
+4, DeePO is activated at $t=6.0\,\mathrm{s}$ and successfully provides real-time control inputs, eliminating the oscillations.
+By comparison, the wind generator system exhibits a low stability margin without DeePO, with persistent oscillations in active and reactive power after $t=6.0\,\mathrm{s}$.
+These sustained oscillations not only threaten the integrity of the wind generator itself, due to potential mechanical resonance, but also compromise the secure operation of the overall power system.
+We observe that the damping performance of DeePO is comparable to that of DeePC (c.f.
 Fig.
-14 in [21]), but the computational burden is significantly less (since DeePC requires solving a quadratic program online).
+14 in [21]).
+Nevertheless, DeePO offers a significant computational advantage.
+For instance, on an AMD Ryzen 7 PRO 7840H CPU with 32 GB of RAM, solving a quadratic program online (required by DeePC) takes approximately $15~{}\mathrm{ms}$ per time step, whereas DeePO only requires around $1~{}\mathrm{ms}$ to perform a gradient step.
+This highlights the efficiency of DeePO in delivering similar control performance at a much lower computational cost.
 Figure 5: The time-domain responses of the wind generator: (a) active power and (b) reactive power. with DeePO controller; with non-adaptive controller.
 In time-varying environments, non-adaptive controllers that rely solely on historical data may fail to maintain effective control.
-In contrast, DeePO adapts by continuously learning from real-time data, allowing it to better accommodate time-varying system conditions.
+To illustrate this, we compare a non-adaptive controller with DeePO under changing system conditions.
+Both controllers apply the same initial policy $K_{t_{0}}$ up to $t=9.0\,\mathrm{s}$.
+At this point, the DC voltage control parameters are reduced to 90% of their nominal values, altering the system dynamics without affecting the equilibrium point.
+After $t=9.0\,\mathrm{s}$, the non-adaptive controller retains the fixed gain $K_{t_{0}}$, while DeePO updates its policy online using real-time input-output data.
 Fig.
-5 shows the time-domain responses of the wind generator using a non-adaptive controller (a fixed controller that achieves similar performance to DeePO before $t=9.0\,\mathrm{s}$) and the DeePO controller, respectively, under parameter changes and disturbances.
-Before $t=9.0\,\mathrm{s}$, both controllers have converged to their respective optimal gains based on previously collected data.
-At $t=9.0\,\mathrm{s}$, the DC voltage control parameters are intentionally reduced to 0.9 times of their original values, simulating a sudden change in the wind generator system.
-While this change does not affect the equilibrium point, it modifies the dynamic characteristics of the system.
-From $t=9.0\,\mathrm{s}$ to $t=21.0\,\mathrm{s}$, the DeePO controller continues to update its control strategy by leveraging real-time data, whereas the non-adaptive controller retains its original control settings without adaptation.
-To ensure that the collected data are informative to update the DeePO controller, noise injection is increased between $t=10.0\,\mathrm{s}$ and $t=17.0\,\mathrm{s}$ to better excite the system and capture its dynamic characteristics.
-At $t=21.0\,\mathrm{s}$, a small disturbance is introduced by injecting a q-axis voltage perturbation into the PLL.
-As shown in Fig.
-5, the DeePO controller quickly damps out oscillations, demonstrating significantly better response performance compared to the non-adaptive controller.
-This improvement can be attributed to the real-time adaptation of the control policy in DeePO, which enables it to perceive changes in system dynamics and maintain superior performance under disturbances.
+5 shows the time-domain responses under this time-varying condition and a subsequent disturbance introduced at $t=21.0\,\mathrm{s}$ by injecting a q-axis voltage perturbation into the PLL.
+While the non-adaptive controller exhibits degraded performance, DeePO successfully maintains stability and quickly damps out oscillations.
+This improvement stems from its ability to continuously adapt to changing dynamics in real time.
 VI Conclusion
-In this paper, we proposed a data-enabled policy optimization (DeePO) method, which is direct data-driven, adaptive, based on online input-output data, and recursive in terms of implementation.
+In this paper, we proposed a output-feedback data-enabled policy optimization (DeePO) method, which is direct data-driven, adaptive, based on online input-output data, and recursive in terms of implementation.
 We demonstrated via simulations the effectiveness of DeePO in stabilizing and mitigating undesired oscillations in a grid-connected power converter and a direct-drive wind generator.
-In particular, we observed that DeePO has satisfactory performance in stabilization, and shows a similar damping ratio to data-enabled predictive control (DeePC).
-Nevertheless, DeePO is computationally more efficient than DeePC, as it is gradient-based and does not require solving optimization problems online.
-Future work includes the extension of DeePO to achieve other control objectives (e.g., reference tracking) and to other types of systems (e.g., slowly time-varying systems).
-Appendix A Proof of Theorem 1
-The state can be represented with system dynamics and history trajectories as
-$x_{t}$ $=A^{n}x_{t-n}+\mathcal{C}{u}_{t,n}+\mathcal{C}_{w}w_{t,n}$ (12)
-${y}_{t,n}$ $=\mathcal{O}x_{t-n}+\mathcal{T}{u}_{t,n}+\mathcal{T}_{w}w_{t,n}+v_{t,n}.$
-Since $\mathcal{O}$ has full column rank, it has a unique left pseudo inverse $\mathcal{O}^{\dagger}=(\mathcal{O}^{\top}\mathcal{O})^{-1}\mathcal{O}^{\top}$.
-Then, it follows immediately from (12) that
-$x_{t}$ $=(\mathcal{C}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}){u}_{t,n}+A^{n}\mathcal{O}^{\dagger}{y}_{t,n}$
-$~{}~{}~{}+(\mathcal{C}_{w}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}_{w})w_{t,n}-A^{n}\mathcal{O}^{\dagger}v_{t,n},$
-and $y_{t}=Cx_{t}+v_{t}=Sz_{t}+d_{t}.$ Thus, a non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization).
-The proof for controllablity of (LABEL:equ:nonnimimal_realization) follows the same vein of that of the key reachability lemma in [30, Lemma 3.4.7].
-Note that their relatively prime condition holds for the system (1) since the length of the past trajectory used to represent the state matches the system order $n$.
+Future work includes the extension of the output-feedback DeePO to achieve other control objectives (e.g., reference tracking) and to other types of systems (e.g., slowly time-varying systems).
+It is also valuable to investigate the effect of regularization on output-feedback DeePO of power converter systems [30].